Steven can afford car payments of $250 a month for 60 months. The bank will lend him this money at 6.0 percent interest. How much can Steven borrow?
Answer #1

This question requires application of PV of annuity formula, according to which

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

n = 60, r = 6%/12 = 0.50% (monthly)

PV = 250 * 1-(1 +0.005) -60 PV = 250* 0.005

PV = 250 * 51.7256

PV = $12,931.39